OpenStudy (vikstar2.0):

Truncate the decimal expansion of 3√6 to find the best approximation. A)1.8 < √6< 1.9 B)1.80 < √6< 1.85 C)1.81 < √6< 1.82 D)1.82 < √6 < 1.83

Parth (parthkohli):

none of those choices are correct. \(\sqrt{6}> \sqrt{4}=2\)

Parth (parthkohli):

Oh, you probably meant:\[\sqrt[3]6\]

OpenStudy (vikstar2.0):

@ParthKohli so whats the answer

OpenStudy (vikstar2.0):

@ParthKohli yea 3√6 is what i meant